# ICP-4D: Bridging Iterative Closest Point and LiDAR Panoptic Segmentation

### v1

@@ C.4 Quantitative Results
We additionally provide the results of each semantic class on panoptic nuScenes [13] to facilitate a more comprehensive understanding. We present our reproduced results obtained from officially given training configurations. Note that 4D-Stop [20] and EQ-4D-Stop [63] do not provide official class-wise results for the panoptic nuScenes benchmark, whereas only overall metrics such as LSTQ are available. For this reason, class-wise comparisons against these methods are excluded to avoid confusion and ensure fairness. In Table A.5, our method yields overall performance improvements, particularly in safety-critical thing classes (e.g., bicycle, car, and pedestrian). Furthermore, classes with relatively fewer points, such as pedestrians and traffic cones, exhibit stronger association performance against the baselines. This observation aligns with our earlier analysis on SemanticKITTI [4] presented in the main manuscript.
C.5 Qualitative Results
@@
In this section, we present additional qualitative analysis by illustrating the results of SemanticKITTI [4] and panoptic nuScenes [13] dataset. As clearly shown in Figure A.2 and  A.3, ICP-4D maintains robust ID consistency even in challenging scenarios. Our method explicitly incorporates the inherent geometric relations between instances and adopts a delicately designed pipeline that is resilient to occlusion and noisy predictions. In contrast, Mask4Former [54] associates instances by selecting the majority ID among spatially overlapped points, making it vulnerable to imprecise predictions and leading to frequent ID switches (see the 1st, 2nd, and 3rd rows of Figure A.2). Apart from this, Mask4D [30] struggles with separating nearby objects and accumulates errors from early mispredictions (see the 1st, 2nd, and 4th rows of Figure A.2). Furthermore, ICP-4D also operates well on the panoptic nuScenes dataset, which is sparse and large temporal gaps between consecutive scans. Specifically, in the 2nd and 4th rows of Figure A.3, we observe that the instances have no overlap and sparse representations. The baselines confuse the instance IDs due to their partial observations and rapid motions. However, ICP-4D successfully finds the corresponding instances even when they are distant. Also, from the 1st row of Figure A.3, we observe that ICP-4D can effectively associate instances in complex scene where pedestrians are heavily entangled.
@@ D.1 Limitation and Future work
We identify several limitations of our proposed approach, which inherits issues from both pre-trained 3D panoptic models and LiDAR sensors. While ICP-4D demonstrates strong association performance with a cost efficient pipeline, real-time deployment remains challenging. Our method highly depends on the performance of the underlying 3D panoptic model; thus, its runtime can be further improved as point cloud processing continues to advance. In addition, partial observation of instances in LiDAR sensing make it difficult to reliably associate instances. Incorporating trajectory cues represents a promising direction to further stabilize instance association. Lastly, our pipeline forces the requirement of camera parameters as a prerequisite for transforming the point coordinates into the world coordinates. Relaxing this requirement would improve the applicability of our method in broader real-world settings. We hope these insights encourage continued progress toward scalable 4D perception systems.
Figure A.2: Qualitative comparison on semanticKITTI [4] validation set. We visualize the association results over five consecutive scans for both the baselines and our ICP-4D. Different colors represent different instances. The dotted boxes zoom in for clear comparison. For clarity, we use circled numbers (①–⑤) to represent the frame indices.
Figure A.3: Qualitative comparison on panoptic nuScenes [13]. We visualize the association results over five consecutive scans for both the baselines and our ICP-4D. Different colors represent different instances. The dotted boxes zoom in for clear comparison. For clarity, we use circled numbers (①–⑤) to represent the frame indices.

### v2

@@ C.5 Quantitative Results
We additionally provide the results of each semantic class on panoptic nuScenes [fong2022panoptic] to facilitate a more comprehensive understanding. We present our reproduced results obtained from officially given training configurations. Note that 4D-Stop [kreuzberg20224d] and EQ-4D-Stop [zhu20234d] do not provide official class-wise results for the panoptic nuScenes benchmark, whereas only overall metrics such as LSTQ are available. For this reason, class-wise comparisons against these methods are excluded to avoid confusion and ensure fairness. In Table A.4, our method yields overall performance improvements, particularly in safety-critical thing classes (e.g., bicycle, car, and pedestrian). Furthermore, classes with relatively fewer points, such as pedestrians and traffic cones, exhibit stronger association performance against the baselines. This observation aligns with our earlier analysis on SemanticKITTI [behley2019semantickitti] presented in the main manuscript.
C.6 Qualitative Results
@@
In this section, we present additional qualitative analysis by illustrating the results of SemanticKITTI [behley2019semantickitti] and panoptic nuScenes [fong2022panoptic] dataset. As clearly shown in Figure A.1 and  A.2, Geo-4D maintains robust ID consistency even in challenging scenarios. Our method explicitly incorporates the inherent geometric relations between instances and adopts a delicately designed pipeline that is resilient to occlusion and noisy predictions. In contrast, Mask4Former [yilmaz2024mask4former] associates instances by selecting the majority ID among spatially overlapped points, making it vulnerable to imprecise predictions and leading to frequent ID switches (see the 1st, 2nd, and 3rd rows of Figure A.1). Apart from this, Mask4D [marcuzzi2023mask4d] struggles with separating nearby objects and accumulates errors from early mispredictions (see the 1st, 2nd, and 4th rows of Figure A.1). Furthermore, Geo-4D also operates well on the panoptic nuScenes dataset, which is sparse and large temporal gaps between consecutive scans. Specifically, in the 2nd and 4th rows of Figure A.2, we observe that the instances have no overlap and sparse representations. The baselines confuse the instance IDs due to their partial observations and rapid motions. However, Geo-4D successfully finds the corresponding instances even when they are distant. Also, from the 1st row of Figure A.2, we observe that Geo-4D can effectively associate instances in complex scene where pedestrians are heavily entangled.
@@ D.1 Limitation and Future work
We identify several limitations of our proposed approach, which inherits issues from both pre-trained 3D panoptic models and LiDAR sensors. While Geo-4D demonstrates strong association performance with a cost efficient pipeline, real-time deployment remains challenging. Our method highly depends on the performance of the underlying 3D panoptic model; thus, its runtime can be further improved as point cloud processing continues to advance. In addition, partial observation of instances in LiDAR sensing make it difficult to reliably associate instances. Incorporating trajectory cues represents a promising direction to further stabilize instance association. Lastly, our pipeline forces the requirement of camera parameters as a prerequisite for transforming the point coordinates into the world coordinates. Relaxing this requirement would improve the applicability of our method in broader real-world settings. We hope these insights encourage continued progress toward scalable 4D perception systems.
Figure A.1: Qualitative comparison on semanticKITTI [behley2019semantickitti] validation set. We visualize the association results over five consecutive scans for both the baselines and our Geo-4D. Different colors represent different instances. The dotted boxes zoom in for clear comparison. For clarity, we use circled numbers (①–⑤) to represent the frame indices.
Figure A.2: Qualitative comparison on panoptic nuScenes [fong2022panoptic]. We visualize the association results over five consecutive scans for both the baselines and our Geo-4D. Different colors represent different instances. The dotted boxes zoom in for clear comparison. For clarity, we use circled numbers (①–⑤) to represent the frame indices.